Circular No. 7873

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1908+094 J. J. M. in ’t Zand, Astronomical Institute, Utrecht University, and

Space Research Organization Netherlands; and M. Capalbi and M. Perri, ASI Science Data Center, Frascati, report on behalf of a larger team that two target-of-opportunity observations were carried out with BeppoSAX on the proposed black-hole x-ray transient XTE J1908+094 (Woods et al., IAUC 7856; Feroci et al., IAUC 7861): “The second observation was with the Medium-Energy Concentrator Spectrometer on Apr. 2.4–3.8 UT and resulted in an order-of-magnitude improvement of the positional accuracy. We localize the source at α = 19^{h}08^{m}52^{s}.6, δ = +9^{o}23′07′′ (equinox 2000.0; uncertainty 17′′, 90-percent confidence). This is 0′.9 from the position determined by Woods et al. but consistent within its 2′ accuracy.”

SUPERNOVA 2002cd IN NGC 6916 T. Matheson, S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002cd (cf. IAUC 7871), obtained by M. Calkins on Apr. 10.49 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type Ia supernova before maximum. Adopting the NASA/IPAC Extragalactic Database recession velocity of 3101 km/s for the host galaxy, the supernova expansion velocity is ∼ 16 800 km/s for Si II (rest 635.5 nm). The spectrum shows strong, narrow Na I D absorptions at zero velocity (0.12-nm equivalent width) and at the velocity of the host galaxy (0.21-nm EW). The dust maps by Schlegel et al. (1998, Ap.J. 500, 525) indicate a Galactic reddening of E(B − V) = 0.41. In addition, the overall shape of the spectrum declines toward the blue, implying that the supernova is heavily extinguished by dust. Aside from the reddening, the spectrum is consistent with a normal Ia about ten days before maximum.

COMETS C/2002 G1 AND C/2002 G2 (SOHO) Further to IAUC 7863, D. Hammer reports his measurements of a non-

Kreutz comet and a Kreutz comet found by R. Kracht on SOHO website C2 coronagraph images. Reduced astrometry and orbital elements by the undersigned appear on the cited MPEC s.

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 G1 Apr. 2.653  0^{h}41.^{m}9  + 3^{o}20′  2002-G13
C/2002 G2  4.579  1 01.4  + 5 08  2002-G28
2002 April 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden